Part B A bill of materials indicates A. what materials the production department needs on the factory floor at that spec...

Part B

A bill of materials indicates

A. what materials the production department needs on the factory floor at that specific moment in time.  

B.what materials are needed to complete the job.  

C.what materials are needed for the entire year.  

D.what materials are needed to be ordered for the month.  

Here are selected data for Whatchamacallit​ Company:

Cost of goods manufactured

$320,600

Work in process​ inventory, beginning

$109,500

Work in process​ inventory, ending

$105,000

Direct materials used

$73,900

Manufacturing overhead is allocated at 60% of direct labor cost.

What was the amount of direct labor​ costs? (Round your answer to the nearest​ cent.)

A.$164,725.00

B.$242,200.00

C.$85,675.00

D.$151,375.00

Widget Maker Manufacturing uses job costing. In​ May, material requisitions were $44,700 ($39,800 of these were direct​ materials), and raw material purchases were $57,700. The end of month balance in raw materials inventory was $25,200.

What was the beginning raw materials inventory​ balance?


A. $12,000

B. $26,800

C. $43,100

D. $72,300

When calculating the total amount of manufacturing overhead to allocate to a particular​ job, the company would multiply each departmental overhead rate by​ ________ and then​ ________ together the allocated amounts from each department.

A. the actual amount of the plantwide allocation based used by the​ job; add

B.the actual amount of the departmental allocation based used by the​ job; multiply

C.the actual amount of the departmental allocation based used by the​ job; add

D.the actual amount of the plantwide allocation based used by the​ job; multiply

Beaver Company manufactures coffee tables and uses an

activityminus−based

costing system to allocate all manufacturing conversion costs. Each coffee tables consists of

3030

separate parts totaling

$ 270$270

in direct​ materials, and requires

77

hours of machine time to produce. Additional information​ follows:

Activity

Allocation Base

Cost Allocation Rate

Materials handling

Number of parts

$2.00

per part

Machining

Machine hours

$3.70

per machine hour

Assembling

Number of parts

$3.00

per part

Packaging

Number of finished units

$7.00

per finished unit

What is the cost of machining per coffee​ table?

A.$25.90

B.$111.00

C.$90.00

D.$49.00

Back Porch Company manufactures lawn chairs using an

activityminus−based

costing system to allocate all manufacturing conversion costs. The company incurs no direct labor costs. The following information is provided for the month of​ June:

Activity

Estimated Indirect

Activity Costs

Allocation Base

Estimated Quantity of

Allocation Base

Materials handling

$17,000

Number of parts

11,200

parts

Assembling

$16,300

Number of parts

11,200

parts

Packaging

$5,500

Number of lawn chairs

1,400

lawn chairs

Each lawn chair consists of 8 ​parts; the total direct materials cost per lawn chair is $16.What is the total manufacturing cost per lawn​ chair? (Round intermediate and the final answer to the nearest​cent.)

A.$39.79

B.$27.71

C.$23.79

D.$ 43.71
Answer #1

1. Option A

A bill of material indicates what materials the production department needs on the factory floor at that specific moment in time.

2. Option D ($151,375)

Calculation of direct labor costs:-

Cost of goods manufactured = Direct materials+Direct labor+Beginning W.I.P-Closing W.I.P+Manufacturing Overhead (MOH)

$320,600 = $73,900 + Direct labor + $109,500 - $105,000 + MOH

$320,600 = Direct labor + $78,400 + MOH

  Direct labor + MOH = $320,600 - $78,400 = $242,200

$242,200 = X + 60% (X)

  $242,200 = X + 0.6 X

1.6 X = $242,200

X = $242,200/ 1.6 = $151,375

Therefore Direct Labor = $151,375

3. Answer is $12,200. No option is matching for the answer. As option is near to the answer, so option A is the answer. But actual answer is $12,200.

Calculation of beginning inventory balance:-

Beginning Inventory Balance = Material requisitions - Raw material purchases + Ending raw material inventory

= $44,700 - $57,700 + $25,200

= $12,200

4. Option C

When calculating the total amount of manufacturing overhead to allocate to a particular​ job, the company would multiply each departmental overhead rate by​ the actual amount of the departmental allocation based used by the​ job and then add together the allocated amounts from each department.

5. Option A ($25.9)

Cost of machining = $3.70 * 7 machine hours

= $25.90

Cost of Material handling = $2 * 30 parts = $60

Cost of Assembling = $3 * 30 parts = $90

Cost of Packaging = $7

Direct materials = $270

Total Manufacturing cost = $452.9

6. Option D ($43.71)

Calculation of total manufacturing cost per lawn chair:-

Material handling = ($17,000 / 11,200 ) * 8 = $12.14

Assembling = ($16,300 / 11,200 ) * 8 = $11.64

Packaging = $5,500 / 1,400 = $3.93

Direct materials = $16

Total Manufacturing cost = $43.71